OpenStudy (anonymous):

Find dy/dx by implicit differentiation. x^5+y^3=4....i got (-5x^4)/(y^2)....

OpenStudy (amistre64):

implicit is just the same as explicit; just a new name thats all

OpenStudy (anonymous):

im just learning the chapter now...

OpenStudy (amistre64):

x^5 + y^3 = 4 Dx(x^5) +Dx(y^3) = Dx(4) x' 5x^4 + y' 3x^2 = 0 ; x'=1; solve for y'

OpenStudy (amistre64):

i typoed it; thats spose to be y' 3y^2 :)

OpenStudy (anonymous):

yes..but wouldnt you have to solve it..to be come..\[3y^2y^'=-5x^4....\]

OpenStudy (anonymous):

and divide it by \[3y^2?\]

OpenStudy (amistre64):

differentiating is the same regardless of what you call it; explicit, implicit or other.... yes, that is correct; the only thing you left out in your post was the '3'

OpenStudy (anonymous):

to get \[-5x^4/3y^2\]

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

ooohh haha silly me

OpenStudy (amistre64):

yeah :) I took a test the other day and derived fine; but I misplaced a sign.. grrr

OpenStudy (anonymous):

i know i hate it when that happens